OpenStudy (anonymous):

A line has slope 2 and (3, 4) is on the line. Which is the correct equation that describes this line in slope-intercept form?

OpenStudy (rulnick):

y = 2x + b but this must work for (3,4), so: 4 = 2(3) + b 4 = 6 + b b = -2 so y = 2x -2 done :)